# Grid-Agent: An LLM-Powered Multi-Agent System for Power Grid Control

### v2

@@ Abstract
Modern power grids face unprecedented complexity from the increasing penetration of Distributed Energy Resources (DERs), the widespread adoption of Electric Vehicles (EVs), and the growing frequency of extreme weather events. This complexity challenges traditional approaches to grid planning, operation, and management. Traditional rule-based systems and numerical optimization approaches often struggle with the scale, dynamics, and adaptability required by modern power networks. This paper introduces Grid-Agent, an autonomous, AI-driven framework that combines Large Language Models (LLMs) within a multi-agent framework to detect and remediate grid violations. Grid-Agent integrates semantic reasoning with numerical precision through a modular agent architecture: a planning agent generates coordinated action sequences using numerical power flow solvers, while a validation agent evaluates system stability and action effectiveness via sandboxed execution with safety rollbacks. To ensure scalability, Grid-Agent incorporates an adaptive multi-scale network representation that dynamically selects optimal encoding schemes based on network size and complexity. The framework enables coordinated violation resolution through optimizing switch configurations, battery deployment, and load curtailment strategies. Experimental results in standard IEEE and CIGRE test systems (IEEE 69-bus, CIGRE MV, and IEEE 30-bus) demonstrate superior violation mitigation performance. Additionally, the framework’s built-in data collection and learning capabilities enable continuous learning and adaptation to diverse network topologies. The autonomous nature of the framework makes it particularly suitable for modern smart grid applications requiring rapid response to dynamic operating conditions.
@@ I Introduction
This paper introduces Grid-Agent, an autonomous, AI-driven framework conceived as an exploration into how LLMs can be applied to power system control, rather than as a finalized, production-ready solution. The primary goal is to push the boundaries of what LLMs and multi-agent systems might bring to future smart grid operations by demonstrating a novel methodology for autonomous grid management. Consequently, this work prioritizes the exploration of agentic reasoning, coordinated planning, and semantic interpretation over the formal optimality of the resulting control actions. The focus is on demonstrating a paradigm where Grid-Agent can interpret network topology, detect electrical violations, and formulate remediation strategies by integrating semantic understanding with domain-specific analysis tools.
The key contributions of this work include:
  • Multi-Agent architecture. We introduce a multi-agent framework that integrates LLM-based planning with numerical validation, enabling autonomous grid violation detection and remediation without human intervention.
  • Adaptive Multi-scale Representation. A dynamic encoding scheme was developed that automatically adjusts network representation granularity based on system size and complexity, ensuring scalability from micro-grids to large distribution networks.
  • Coordinated Action Optimization. A coordinated violation resolution framework that optimizes multi-violation solutions through topology reconfiguration, battery deployment, and demand response that minimize the number of control actions while maximizing violation resolution effectiveness.
  • Comprehensive Experimental Validation. The framework’s performance was validated on standard IEEE and CIGRE test systems, showing superior violation mitigation.
  • Continuous Learning Capability. A training data generation framework that creates high-quality datasets in multiple formats for LLM fine-tuning to enable the system to learn from operational experience, ensuring its performance improves and adapts over time.
@@ II-A Power Systems Optimization
Traditional power system optimization has progressed from manual control to sophisticated automated frameworks. Classical methods, such as Optimal Power Flow (OPF) [4], [5], have been widely applied to objectives like peak shaving, cost/loss minimization, and violation mitigation. While mathematically robust, these approaches often face convergence issues due to the complexity of modern grids, which incorporate high levels of distributed renewable energy, highly variable loads, and numerous discrete controllable assets such as switches and tap changers. These factors result in a mixed-integer, non-convex, nonlinear, and unbalanced AC optimal powerflow problem, which is NP-hard and computationally intensive to solve. Furthermore, their rigid mathematical structure makes it difficult to incorporate heuristic operator knowledge or adapt to unforeseen event types. Furthermore, these approaches often address isolated tasks and lack a unified framework for holistic, system-level optimization and control. This highlights the need for a system that can integrate high-level strategic reasoning with precise numerical validation, a core principle of our proposed Grid-Agent framework.
Recent advances in artificial intelligence (AI) have brought machine learning techniques into power systems, with promising results in tasks such as load forecasting [6], renewable generation prediction [7], and fault detection [8]. However, these approaches generally address isolated tasks and lack a unified framework for holistic, system-level optimization and control.
@@ II-B Large Language Models in Engineering Applications
The application of LLMs to engineering domains has gained significant attention following the success of models like GPT-4 and Claude. Recent work has explored LLM applications in software engineering [9], materials science [10], and system design [11].
In the energy sector, preliminary investigations have examined LLMs for grid analytics [12], energy trading [13], and demand response optimization [14]. However, these applications primarily focus on data analysis and prediction rather than direct system control and optimization. This leaves a critical gap in utilizing LLM reasoning for active, real-time control and optimization in critical infrastructure, which is the primary focus of our work.
@@ II-C Multi-Agent Systems with LLMs
The integration of LLMs with multi-agent systems is an emerging and rapidly advancing area of research. Recent studies have demonstrated the potential of LLM-based multi-agent frameworks in domains such as software engineering [15], automated scientific research [16], and collaborative problem-solving [17]. These systems leverage the reasoning capabilities of LLMs while maintaining structured workflows and state management.
However, the application of LLM-based multi-agent systems to critical infrastructure domains like power grids presents unique challenges including safety requirements, real-time constraints, and the need for explainable decisions in mission-critical environments.
III Preliminaries
@@ III-A Power Grid Violation Resolution Problem
We model the power network as a graph G=(N,E){G}=({N},{E}), where ${N}=\{1,2,\dots,n\}$ is the set of buses (nodes) and E{E} is the set of transmission lines and transformers (edges). The state of the grid is defined by the voltage magnitude $V_{i}$ and angle $\delta_{i}$ at each bus $i\in{N}$, along with the active power $P_{ij}$, reactive power $Q_{ij}$, and current flow $I_{ij}$ for each line $(i,j)\in{E}$.
A violation $\mathcal{V}$ occurs when the system state breaches predefined operational limits. We consider three primary categories of violations: voltage violations, thermal violations, and disconnected buses.
A voltage violation $\mathcal{V}_{\text{volt}}$ occurs when a bus voltage magnitude $V_{i}$ falls outside its acceptable range $V_{\min},V_{\max}]$. Formally, the set of buses experiencing voltage violations is defined as
$\mathcal{V}_{\text{volt}}=\{i\in N\mid V_{i}<V_{\min}\text{ or }V_{i}>V_{\max}\}.$ (1)
A thermal violation $\mathcal{V}_{\text{therm}}$ arises when the apparent power flow $S_{ij}$ or current $I_{ij}$ on a line exceeds its maximum rating. This can be expressed as
$\mathcal{V}_{\text{therm}}=\{(i,j)\in E\mid S_{ij}>S_{ij}^{\max}\text{ or }I_{ij}>I_{ij}^{\max}\}.$ (2)
@@ III-C Action Space Constraints
The control actions available to Grid-Agent are constrained by the physical and operational capabilities of the grid equipment. The action space consists of switch operations, load curtailment, and battery placement with dispatch.
Switch operations are represented as binary decisions to open or close a switchable line $(i,j)$ within the set of designated switchable lines ${E}_{\text{sw}}\subseteq{E}$. For each switchable line, the state variable is defined as
@@
Load curtailment is modeled as a continuous reduction in the active power demand $P_{i}^{\text{load}}$ at buses belonging to the set of curtailable loads ${N}_{\text{L}}\subseteq{N}$. The reduction is parameterized by a curtailment fraction $\gamma_{i}$, resulting in a new load value given by
$P_{i}^{\text{new\_load}}=P_{i}^{\text{orig\_load}}\cdot(1-\gamma_{i}),$ (6)
$0\leq\gamma_{i}\leq\gamma_{\max},\quad\forall i\in{N}_{\text{L}}.$ (7)
Battery placement and dispatch involve two coupled decisions: a discrete choice of where to place batteries, and a continuous choice of how much active and reactive power to inject or absorb. Placement is modeled as a binary decision variable $b_{i}$ for each bus, constrained by the total number of available batteries $B_{\text{total}}$:
@@
$P_{B,i}^{2}+Q_{B,i}^{2}\leq(b_{i}\cdot S_{B}^{\max})^{2},$ (9)
$-P_{B}^{\max}\leq P_{B,i}\leq P_{B}^{\max},$ (10)
$-Q_{B}^{\max}\leq Q_{B,i}\leq Q_{B}^{\max}.$ (11)
This formulation guarantees that batteries are both allocated and dispatched within their physical limits, ensuring they can effectively support grid operations.
@@ III-D2 Multi-Agent Systems for Decomposed Problem Solving
The multi-agent system can be formulated as
@@
IV System Architecture and Methodology
@@ IV-A State-Driven Multi-Agent Workflow
Figure 1: The architecture diagram of the Grid Agent.
@@ IV-B LLM-Guided Coordinated Action Planning
The central task of finding an optimal action sequence is navigated by the Planner Agent. Unlike traditional optimizers, Grid-Agent uses the LLM as a tool-augmented reasoning engine. This is achieved through a comprehensive system prompt that structures the LLM’s task, defining five key elements:
The core of this process is the Planner Agent, which formulates a multi-step plan (Algorithm 1).The agent’s reasoning is guided by a structured system prompt that specifies:
1) Role Definition: The LLM is instructed to act as an expert power system operator, grounding its responses in domain-specific best practices.
2) State Context: It receives a structured representation of the current network state, including topology, component statuses, and a detailed list of all active violations.
3) Action Space and Constraints: The prompt explicitly defines the set of available actions (e.g., update_switch_status, add_battery, curtail_load) and their associated operational constraints (e.g., maximum number of deployable batteries, available curtailable loads).
4) Strategic Guidance: The LLM is guided by a prioritized action policy that mirrors industry best practices: (1) Topology Reconfiguration, (2) Battery Deployment/Dispatch, and (3) Load Curtailment. This constrains the search space and aligns the LLM’s reasoning with established operational logic.
5) Output Schema: To ensure reliable machine integration, the LLM is required to respond exclusively with a machine-parsable list of tool calls in a structured format (e.g., JSON). This eliminates ambiguous natural language and enables direct interfacing with the Executor Agent.
@@ IV-C Adaptive Multi-Scale Network Representation
A primary challenge in applying LLMs to large-scale networks is the constraint of the model’s context window. To overcome this, Grid-Agent employs an adaptive network representation scheme that dynamically balances information fidelity with computational tractability. The framework automatically selects the level of detail based on network size and complexity:
Full-Component Detail: For smaller networks, the system provides a complete serialization of every component (buses, lines, loads, generators), offering the LLM comprehensive data for fine-grained analysis.
Semantic Graph Abstraction: For larger networks or scenarios with clustered violations, the system generates a semantic graph representation. This abstraction summarizes healthy sections of the grid and focuses on the electrical relationships between violated components and nearby controllable assets, enabling the LLM to identify coordinated, system-level solutions efficiently.
This multi-scale approach ensures the LLM receives contextually relevant information without being overwhelmed by excessive data, enabling the framework to scale from microgrids to large distribution networks.
@@ V-A Test Network Configuration
We validated Grid-Agent’s performance on a diverse set of standard IEEE and CIGRE test systems to ensure applicability across networks of varying size, topology, and complexity. The specific test cases were generated by introducing multiple, often interacting, violations to stress-test the agent’s reasoning and coordinated control capabilities.
For the IEEE 30-Bus System, two transmission scenarios were created: Case30 Light, a baseline with 2 voltage and 3 thermal violations, and Case30 Medium, a simpler thermal violation scenario.
To simulate severe distribution-level challenges, we used the CIGRE Medium Voltage (MV) Distribution Network, a 15-bus European benchmark, featuring the CIGRE MV Severe scenario with 14 concurrent violations and the CIGRE MV Disconnected scenario, where a fault isolates 5 buses.
Finally, to evaluate scalability, the larger IEEE 69-Bus Distribution System was tested with scenarios including IEEE 69 Large Loads (18 simultaneous violations), IEEE 69 Medium Loads (a highly complex case with 29 violations), and IEEE 69 Disconnected (a fault resulting in 13 disconnected buses).
An automated script was used to generate these violation scenarios, ensuring repeatable and comprehensive testing across different violation types and severity levels. Table I summarizes the specifications of the benchmark networks employed in this study.
@@ V-B LLM Test Suite Configuration
Parameter IEEE Case30 CIGRE MV IEEE 69-bus
@@
Voltage Level 132-33 kV 20 kV 12.66 kV
TABLE II: Large Language Model Configurations for Benchmarking
@@ V-C Performance Metrics
System performance was evaluated using a suite of quantitative metrics designed to capture the effectiveness, efficiency, and quality of the violation resolution process. The success rate served as the primary metric, defined as the percentage of violation scenarios that were successfully and fully resolved.
To assess the efficiency of control actions, we measured action efficiency, defined as the average number of violations resolved per control action. This metric reflects the agent’s ability to formulate coordinated and impactful solutions rather than relying on redundant or scattered decisions. Another important aspect was convergence speed, which captures the average number of planning–execution–validation iterations required to achieve complete resolution.
Beyond efficiency and speed, we also considered solution quality. This metric emphasizes not only the resolution of violations but also the manner in which solutions are achieved. High-quality solutions minimize the number of control actions and avoid disruptive measures, such as load curtailment, whenever possible. Finally, runtime was recorded as the average wall-clock time, measured in seconds, required to reach a final solution.
Together, these metrics provide a comprehensive evaluation of the system’s performance, balancing effectiveness, efficiency, practicality, and solution robustness.

### v3

@@ Abstract
Modern power grids face unprecedented complexity from Distributed Energy Resources (DERs), Electric Vehicles (EVs), and extreme weather, while also being increasingly exposed to cyberattacks that can trigger grid violations. This paper introduces Grid-Agent, an autonomous AI-driven framework that leverages Large Language Models (LLMs) within a multi-agent system to detect and remediate violations. Grid-Agent integrates semantic reasoning with numerical precision through modular agents: a planning agent generates coordinated action sequences using power flow solvers, while a validation agent ensures stability and safety through sandboxed execution with rollback mechanisms. To enhance scalability, the framework employs an adaptive multi-scale network representation that dynamically adjusts encoding schemes based on system size and complexity. Violation resolution is achieved through optimizing switch configurations, battery deployment, and load curtailment. Our experiments on IEEE and CIGRE benchmark networks, including the IEEE 69-bus, CIGRE MV, IEEE 30-bus test systems, demonstrate superior mitigation performance, highlighting Grid-Agent’s suitability for modern smart grids requiring rapid, adaptive response.
@@ I Introduction
This paper introduces Grid-Agent, an autonomous, AI-driven framework conceived as an exploration into how LLMs can be applied to power system control, rather than as a finalized, production-ready solution. The primary goal is to push the boundaries of what LLMs and multi-agent systems might bring to future smart grid operations by demonstrating a novel methodology for autonomous grid management. Consequently, this work prioritizes the exploration of agentic reasoning, coordinated planning, and semantic interpretation over the formal optimality of the resulting control actions. The focus is on demonstrating a paradigm where Grid-Agent can interpret network topology, detect electrical violations, and formulate remediation strategies by integrating semantic understanding with domain-specific analysis tools. The key contribution of this work includes:
  • Introduction of Grid-Agent, a modular multi-agent framework that combines LLM-based semantic reasoning with numerical solvers to autonomously detect and remediate grid violations. To ensure scalability across diverse network sizes, Grid-Agent incorporates an adaptive multi-scale representation that dynamically adjusts encoding granularity. Violation resolution is formulated as a coordinated optimization problem that integrates topology reconfiguration, battery scheduling, and demand response to maximize effectiveness while minimizing the number of control actions. Furthermore, the framework embeds a continuous learning pipeline that generates training data and fine-tunes its LLM components, enabling performance improvement and adaptation to evolving grid conditions over time.
  • To validate the above contribution, Grid-Agent was implemented and tested on benchmark IEEE and CIGRE systems, including the IEEE 69-bus, the IEEE 30-bus, and the CIGRE MV test systems. Our experimental results demonstrate its superior ability to mitigate violations compared with baseline approaches, its scalability across networks of different sizes and complexities, and its robustness against multiple simultaneous violations, including those arising from cyberattack-induced contingencies.
@@ II-A Power Systems Optimization
Traditional power system optimization has progressed from manual control to sophisticated automated frameworks. Classical methods, such as Optimal Power Flow (OPF) [4], [5], have been widely applied to objectives like peak shaving, cost/loss minimization, and violation mitigation. While mathematically robust, these approaches often face convergence issues due to the complexity of modern grids, which incorporate high levels of distributed renewable energy, highly variable loads, and numerous discrete controllable assets such as switches and tap changers. These factors result in a mixed-integer, non-convex, nonlinear, and unbalanced AC optimal powerflow problem, which is NP-hard and computationally intensive to solve. Furthermore, their rigid mathematical structure makes it difficult to incorporate heuristic operator knowledge or adapt to unforeseen event types. Furthermore, these approaches often address isolated tasks and lack a unified framework for holistic, system-level optimization and control. This highlights the need for a system that can integrate high-level strategic reasoning with precise numerical validation, a core principle of our proposed Grid-Agent framework. Recent advances in artificial intelligence (AI) have brought machine learning techniques into power systems, with promising results in tasks such as load forecasting [6], renewable generation prediction [7], and fault detection [8]. However, these approaches generally address isolated tasks and lack a unified framework for holistic, system-level optimization and control.
@@ II-B Large Language Models in Engineering Applications
The application of LLMs to engineering domains has gained significant attention following the success of models like GPT-4 and Claude. Recent work has explored LLM applications in software engineering [9], materials science [10], and system design [11]. In the energy sector, preliminary investigations have examined LLMs for grid analytics [12], energy trading [13], and demand response optimization [14]. However, these applications primarily focus on data analysis and prediction rather than direct system control and optimization. This leaves a critical gap in utilizing LLM reasoning for active, real-time control and optimization in critical infrastructure, which is the primary focus of our work.
@@ II-C Multi-Agent Systems with LLMs
The integration of LLMs with multi-agent systems is an emerging and rapidly advancing area of research. Recent studies have demonstrated the potential of LLM-based multi-agent frameworks in domains such as software engineering [15], automated scientific research [16], and collaborative problem-solving [17]. These systems leverage the reasoning capabilities of LLMs while maintaining structured workflows and state management. However, the application of LLM-based multi-agent systems to critical infrastructure domains like power grids presents unique challenges including safety requirements, real-time constraints, and the need for explainable decisions in mission-critical environments.
III Grid-Agent: Preliminaries
@@ III-A Power Grid Violation Resolution Problem
We model the power network as a graph G=(N,E){G}=({N},{E}), where ${N}=\{1,2,\dots,n\}$ is the set of buses (nodes) and E{E} is the set of transmission lines and transformers (edges). The state of the grid is defined by the voltage magnitude $V_{i}$ and angle $\delta_{i}$ at each bus $i\in{N}$, along with the active power $P_{ij}$, reactive power $Q_{ij}$, and current flow $I_{ij}$ for each line $(i,j)\in{E}$. A violation $\mathcal{V}$ occurs when the system state breaches predefined operational limits. We consider three primary categories of violations: voltage violations, thermal violations, and disconnected buses. A voltage violation $\mathcal{V}_{\text{volt}}$ occurs when a bus voltage magnitude $V_{i}$ falls outside its acceptable range $V_{\min},V_{\max}]$. Formally, the set of buses experiencing voltage violations is defined as
$\mathcal{V}_{\text{volt}}=\{i\in N\mid V_{i}<V_{\min}\text{ or }V_{i}>V_{\max}\}$ (1)
Moreover, a thermal violation $\mathcal{V}_{\text{therm}}$ arises when the apparent power flow $S_{ij}$ or current $I_{ij}$ on a line exceeds its maximum rating. This can be expressed as
$\mathcal{V}_{\text{therm}}=\{(i,j)\in E\mid S_{ij}>S_{ij}^{\max}\text{ or }I_{ij}>I_{ij}^{\max}\}$ (2)
@@ III-C Action Space Constraints
The control actions available to Grid-Agent are constrained by the physical and operational capabilities of the grid equipment. The action space consists of switch operations, load curtailment, and battery placement with dispatch. Switch operations are represented as binary decisions to open or close a switchable line $(i,j)$ within the set of designated switchable lines ${E}_{\text{sw}}\subseteq{E}$. For each switchable line, the state variable is defined as
@@
Herein, load curtailment is modeled as a continuous reduction in the active power demand $P_{i}^{\text{load}}$ at buses belonging to the set of curtailable loads ${N}_{\text{L}}\subseteq{N}$. The reduction is parameterized by a curtailment fraction $\gamma_{i}$, resulting in a new load value given by
$P_{i}^{\text{new\_load}}=P_{i}^{\text{orig\_load}}\cdot(1-\gamma_{i})$ (6)
$0\leq\gamma_{i}\leq\gamma_{\max},\quad\forall i\in N_{\text{L}}$ (7)
Further, battery placement and dispatch involve two coupled decisions: a discrete choice of where to place batteries, and a continuous choice of how much active and reactive power to inject or absorb. Placement is modeled as a binary decision variable $b_{i}$ for each bus, constrained by the total number of available batteries $B_{\text{total}}$:
@@
$P_{B,i}^{2}+Q_{B,i}^{2}\leq(b_{i}\cdot S_{B}^{\max})^{2}$ (9)
$P_{B}^{\max}\leq P_{B,i}\leq P_{B}^{\max}$ (10)
$Q_{B}^{\max}\leq Q_{B,i}\leq Q_{B}^{\max}$ (11)
The above formulation guarantees that batteries are both allocated and dispatched within their physical limits, ensuring they can effectively support grid operations.
@@ III-D2 Multi-Agent Systems for Decomposed Problem Solving
Based on the above, the multi-agent system can be formulated as
@@
IV Grid-Agent: Proposed System Architecture and Methodology
@@ IV-A State-Driven Multi-Agent Workflow
Figure 1: Conceptual diagram of the Proposed Grid Agent
@@ IV-B LLM-Guided Coordinated Action Planning
The central task of finding an optimal action sequence is navigated by the Planner Agent. Unlike traditional optimizers, Grid-Agent uses the LLM as a tool-augmented reasoning engine. This is achieved through a comprehensive system prompt that structures the LLM’s task, defining five key elements: The core of this process is the Planner Agent, which formulates a multi-step plan (Algorithm 1).The agent’s reasoning is guided by a structured system prompt that specifies: 1) Role Definition: The LLM is instructed to act as an expert power system operator, grounding its responses in domain-specific best practices. 2) State Context: It receives a structured representation of the current network state, including topology, component statuses, and a detailed list of all active violations. 3) Action Space and Constraints: The prompt explicitly defines the set of available actions (e.g., update_switch_status, add_battery, curtail_load) and their associated operational constraints (e.g., maximum number of deployable batteries, available curtailable loads). 4) Strategic Guidance: The LLM is guided by a prioritized action policy that mirrors industry best practices: (1) Topology Reconfiguration, (2) Battery Deployment/Dispatch, and (3) Load Curtailment. This constrains the search space and aligns the LLM’s reasoning with established operational logic. 5) Output Schema: To ensure reliable machine integration, the LLM is required to respond exclusively with a machine-parsable list of tool calls in a structured format (e.g., JSON). This eliminates ambiguous natural language and enables direct interfacing with the Executor Agent.
@@ IV-C Adaptive Multi-Scale Network Representation
A primary challenge in applying LLMs to large-scale networks is the constraint of the model’s context window. To overcome this, Grid-Agent employs an adaptive network representation scheme that dynamically balances information fidelity with computational tractability. The framework automatically selects the level of detail based on network size and complexity. This is performed on two scales: 1) Full-Component Detail and 2) Semantic Graph Abstraction. On the full-component detail scale, for smaller networks, the system provides a complete serialization of every component (buses, lines, loads, generators), offering the LLM comprehensive data for fine-grained analysis. On the semantic-graph abstraction scale, for larger networks or scenarios with clustered violations, the system generates a semantic graph representation. This abstraction summarizes healthy sections of the grid and focuses on the electrical relationships between violated components and nearby controllable assets, enabling the LLM to identify coordinated, system-level solutions efficiently. This multi-scale approach ensures the LLM receives contextually relevant information without being overwhelmed by excessive data, enabling the framework to scale from microgrids to large distribution networks.
@@ V-A Test Network Configuration
We validated Grid-Agent’s performance on a diverse set of standard IEEE and CIGRE test systems to ensure applicability across networks of varying size, topology, and complexity. The specific test cases were generated by introducing multiple, often interacting, violations to stress-test the agent’s reasoning and coordinated control capabilities. For the IEEE 30-Bus System, two transmission scenarios were created: Case30 Light, a baseline with 2 voltage and 3 thermal violations, and Case30 Medium, a simpler thermal violation scenario. To simulate severe distribution-level challenges, we used the CIGRE Medium Voltage (MV) Distribution Network, a 15-bus European benchmark, featuring the CIGRE MV Severe scenario with 14 concurrent violations and the CIGRE MV Disconnected scenario, where a fault isolates 5 buses. Finally, to evaluate scalability, the larger IEEE 69-Bus Distribution System was tested with scenarios including IEEE 69 Large Loads (18 simultaneous violations), IEEE 69 Medium Loads (a highly complex case with 29 violations), and IEEE 69 Disconnected (a fault resulting in 13 disconnected buses). An automated script was used to generate these violation scenarios, ensuring repeatable and comprehensive testing across different violation types and severity levels. Table I summarizes the specifications of the benchmark networks employed in this study.
@@ V-B LLM Test Suite Configuration
  Parameter System IEEE Case30 CIGRE MV IEEE 69-bus
@@
Voltage Level 132–33 kV 20 kV 12.66 kV
TABLE II: Investigated LLM Configurations for Benchmarking
@@ V-C Performance Metrics
System performance was evaluated using a suite of quantitative metrics designed to capture the effectiveness, efficiency, and quality of the violation resolution process. The success rate served as the primary metric, defined as the percentage of violation scenarios that were successfully and fully resolved. To assess the efficiency of control actions, we measured action efficiency, defined as the average number of violations resolved per control action. This metric reflects the agent’s ability to formulate coordinated and impactful solutions rather than relying on redundant or scattered decisions. Another important aspect was convergence speed, which captures the average number of planning–execution–validation iterations required to achieve complete resolution. Beyond efficiency and speed, we also considered solution quality. This metric emphasizes not only the resolution of violations but also the manner in which solutions are achieved. High-quality solutions minimize the number of control actions and avoid disruptive measures, such as load curtailment, whenever possible. Finally, runtime was recorded as the average wall-clock time, measured in seconds, required to reach a final solution. Together, these metrics provide a comprehensive evaluation of the system’s performance, balancing effectiveness, efficiency, practicality, and solution robustness.